Check the continuity of the function f: R^2 to R at (0,0)

f(x,y) = { 3x^2y/(x^2+y^2) if (x,y)≠(0,0)

{ 3 if (x,y)= (0,0)


1
Expert's answer
2021-06-22T11:16:20-0400

Approaching "(0,0)" along the line "y=x"


"\lim\limits_{(x,y)\to(0,0)}\dfrac{3x^2y}{x^2+y^2}=\lim\limits_{(x,y)\to(0,0)}\dfrac{3x^3}{2x^2}=0"

Since "f(0, 0)=3\not=0," the function "f" is not continous at "(0, 0)."